Mrs. Donovan is the coach of a middle school golf team that has 10 golfers. There are 2 sixth graders, 4 seventh graders, and 4 eighth graders on the team. Mrs. Donovan will randomly select two team captains.
What is the probability that both captains will be eighth graders?

Answer by edjones(8007)

2/10 * 1/9 = 2/90 = 1/45
.
Ed